OpenStudy (anonymous):

A certain forest covers an area of 4500km2 . Suppose that each year this area decreases by 8% . What will the area be after 9 years? Use the calculator provided and round your answer to the nearest square kilometer.

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (bloomlocke367):

well, it decreases by 8% in one year.. you need to find out how much it decreases in 9 years.. so you need to make a proportion

OpenStudy (anonymous):

the explanation has like an entire page of steps to take to get the answer but I feel like there is a simpler way to write the equation =o

OpenStudy (anonymous):

I have something along the lines of 4500 - (8% *9)

OpenStudy (bloomlocke367):

ehhh, not quite.

OpenStudy (bloomlocke367):

let's just find the percent for 9 years first

OpenStudy (anonymous):

which would be 0.08 * 9 right?

OpenStudy (bloomlocke367):

yes

OpenStudy (anonymous):

mmk that's 0.72

OpenStudy (bloomlocke367):

now you have to find out what is 72% of 4500

OpenStudy (anonymous):

3240

OpenStudy (bloomlocke367):

yes, now subtract that from your initial amount

OpenStudy (anonymous):

1260

OpenStudy (bloomlocke367):

yepyep!

OpenStudy (bloomlocke367):

oops XD I only meant for that to be there once

OpenStudy (anonymous):

lol it's ok thank you =)

OpenStudy (bloomlocke367):

you're welcome!

OpenStudy (anonymous):

they say the answer is actually 2125 =/

OpenStudy (bloomlocke367):

how? hang on..

OpenStudy (anonymous):

no idea the explanation is long and a little all together I prefer my explanations spread out in like a organized step by step kinda format hahaha

OpenStudy (bloomlocke367):

ummm can you put it here?.. I'll read over it and find the error and explain it to you

OpenStudy (anonymous):

hold on I'll try

OpenStudy (anonymous):

oh well it's easier if I post it then edit it because it messes up the equations and what not when I copy and paste

OpenStudy (anonymous):

or not it's to long.. maybe I'll just type out the equation in order of how they put it and see what you think

OpenStudy (bloomlocke367):

okay

OpenStudy (anonymous):

it starts off talking about the percent of the previous years area. 4500 * 0.92 (4500 * 0.92) * 0.92 = 4500 * 0.92^2 (4500 * 0.92^2) * 0.92 = 4500 * 0.92^3 4500 * 0.92^t 4500 * 0.92^9 = 2124.72 2125

OpenStudy (bloomlocke367):

where did the 0.92 come from?

OpenStudy (anonymous):

The area that the forest covers decreases by 8% each year. This means that the area is 92% of the previous year's area.

OpenStudy (anonymous):

I'm not entirely sure how that works

OpenStudy (bloomlocke367):

ahhhh that makes sense.

OpenStudy (bloomlocke367):

I'm sorry for helping you in the wrong way

OpenStudy (anonymous):

i'm glad because i don't get it lol

OpenStudy (anonymous):

and oh no ! it's ok! lo;

OpenStudy (anonymous):

lol* I'm just grateful for the help

OpenStudy (bloomlocke367):

okay... well I'm not sure how to explain what the explanation means.. sorry :/

OpenStudy (anonymous):

it's ok =) thank you anyways! =)

OpenStudy (bloomlocke367):

you're welcome :)